# Enabling stratified sampling in high dimensions via nonlinear dimensionality reduction

### v1

@@ 4.2 Performance of the heuristic algorithm
Figure 6: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the proposed estimator $\widehat{q}_{\mathrm{sMC}}$ (solid line), for different allocation strategies and stratification approaches: uniform (u), halved (h), optimal (o). The gray dash-dotted vertical line represents the exact value of the quantity of interest.
We now investigate possible improvements achieved by the heuristic algorithm introduced in Section 2.3. We fix $M=100$, $K=10^{6}$, $S=10$, assume a computational budget $N=1,000$, and test two allocation strategies, as well as two alternatives to iteratively refine the strata. The comparison with standard Monte Carlo and Monte Carlo with Cartesian stratification is shown in Figure 6. We observe that the heuristic algorithm is always able to outperform uniform stratification, but we do not notice a significant difference between following a bisection approach, i.e., dividing in two equal parts the interval that contributes the most to the final variance, and selecting the optimal new stratum at each step. Therefore, in order to limit the additional computational overhead, we suggest to use the former approach. We finally remark that, as predicted by theory (see equation (2.20)), the optimal allocation (1) gives a variance which is always smaller than the one produced by a proportional allocation (2).
@@ 4.3 Comparison with active subspaces
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and stratified Monte Carlo estimator $\widehat{q}_{\mathrm{sMC}}$ (solid line) with both AS and NeurAM-based stratification, varying the number of strata $S=2$ (left) and $S=8$ (right). The gray dash-dotted vertical line represents the exact value of the quantity of interest. Top: stratification of the domain. Bottom: distribution of the estimators.
NeurAM has the advantages of being nonlinear and entirely data-driven, requiring no knowledge of the model gradient, but it is not the only available option to improve stratified sampling. In this section, we demonstrate that stratified sampling can be combined with other dimensionality reduction techniques. However, we also show that NeurAM outperforms linear state-of-the-art approaches such as the active subspace method [2]. Inspired by [29], we build a map from the original domain to the unit interval as follows. Let $\mathcal{G}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a transformation that maps a standard Gaussian $\mathcal{N}(0,I)$ into the input distribution $\mu$, i.e. $\mathcal{G}_{\#}\mathcal{N}(0,I)=\mu$, and define the reparameterized model
@@
where the function $\operatorname{erf}$ is evaluated componentwise, otherwise the map can be computed using, e.g., normalizing flows [16]. Then, let $B$ be the matrix
@@
which is symmetric positive semidefinite and therefore has positive eigenvalues. We denote by $v$ the eigenvector corresponding to the largest eigenvalue, representing the one-dimensional active subspace, and normalize it to have unit norm. We note that the map $\mathcal{G}$ is important because Gaussian distributions are preserved under linear transformations and, in particular, if $z\sim\mathcal{N}(0,I)$ then $v^{\top}z\sim\mathcal{N}(0,1)$. Therefore, a point $x\in\mathbb{D}$ can be mapped into the unit interval $[0,1]$ through the one-dimensional AS, by applying the inverse CDF of the standard Gaussian distribution to $v^{\top}\mathcal{G}^{-1}(x)$. Specifically, using the notation of Section 2.1, equation (2.12) now reads
@@
In Figure 7 we compare the stratification provided by AS with the NeurAM-based stratification for a uniform division of the unit interval into $S=2$ and $S=8$ strata, and a computational budget of $N=1024$. The distribution of the corresponding stratified Monte Carlo estimators with proportional allocation is also shown and compared to that of standard Monte Carlo, which is consistently outperformed. While the results are comparable for $S=2$, we observe that the estimator based on nonlinear dimensionality reduction achieves significantly greater variance reduction than linear techniques when the number of strata increases, e.g., $S=8$. This improvement is primarily due to the greater expressiveness of nonlinear methods, which are better able to follow the contour lines of nonlinear models.
@@ 4.4 Stratified multifidelity estimators
Figure 8: Comparison between standard (multifidelity) Monte Carlo estimators $\widehat{q}_{\mathrm{MC}},\widehat{q}_{\mathrm{MFMC}}$ (dashed line) and the proposed estimators $\widehat{q}_{\mathrm{sMC}},\widehat{q}_{\mathrm{sMFMC}}$ (solid line), for different allocation strategies. The gray dash-dotted vertical line represents the exact value of the quantity of interest.
In this section, we test the performance of estimators combining NeurAM stratification and multifidelity variance reduction, as discussed in Section 3. Let $\mathcal{Q}^{\mathrm{HF}}=\mathcal{Q}_{0}$, and consider a low-fidelity model of the form
@@
for which we assume a cost ratio $w=0.01$, to reflect cost differences in realistic applications. This poorly correlated low-fidelity model has already been used in [5, 30, 29, 28]. Following [28, Section 3.2], we then train NeurAM for both models using $M=100$, and re-parameterize the low-fidelity model using
@@
resulting in higher correlations with respect to the high-fidelity model. In Figure 8 we compare (stratified) Monte Carlo estimators with (stratified) multifidelity Monte Carlo estimators for both the allocation strategies, assuming a computational budget $N=2,000$ and using a uniform stratification with $S=5$ strata. First, we observe that stratification is always beneficial, and stratified estimators outperform multifidelity estimators. Then, we notice that leveraging stratification for multifidelity estimators allows us to obtain a significant variance reduction. Finally, we remark that, when using the optimal allocation strategy (1), the improvement with respect to single-fidelity estimators is greater. This is in agreement with Remarks 3.3 and 3.2, where a weaker sufficient condition is provided for allocation (1) to achieve variance reduction.
4.5 High-dimensional models
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the stratified estimators $\widehat{q}_{\mathrm{sMC}}$ (solid line) with both standard (if possible) and NeurAM-based stratification, for the four models in Section 4.5.
After investigating the convergence of the proposed approach and its sensitivity to hyperparameters, we now focus on how these properties scale to high dimensions. Consider the following models:
  • $\mathcal{Q}_{1}\colon\mathbb{R}^{3}\to\mathbb{R}$ proposed in [14] which exhibits strong nonlinearity and nonmonotonicity
@@ 4.5 High-dimensional models
  • $\mathcal{Q}_{2}\colon\mathbb{R}^{4}\to\mathbb{R}$ employed in [8] which models the average velocity of a steady, incompressible, and laminar flow of an electrically conducting fluid between two infinite parallel plates in the presence of a uniform magnetic field (see Hartmann problem, e.g. [25])
@@
  • $\mathcal{Q}_{3}\colon\mathbb{R}^{8}\to\mathbb{R}$ proposed in [12] as a model for the flow of water through a borehole
@@
    $\mathcal{Q}_{4}(x)=\prod_{i=1}^{10}\frac{\left\lvert x_{i}-1\right\rvert+i}{1+i}.$ (4.12)
@@
    $\mu_{3}$ $=\mathcal{N}(0.10,0.0161812^{2})\otimes\log\mathcal{N}(7.71,1.0056^{2})\otimes\mathcal{U}([63070,115600])\otimes\mathcal{U}([990,1110])$ (4.13)
    $\quad\otimes\mathcal{U}([63.1,116])\otimes\mathcal{U}([700,820])\otimes\mathcal{U}([1120,1680])\otimes\mathcal{U}([9855,12045]);$
@@
where $\log\mathcal{U}$ and $\log\mathcal{N}$ denote the log-uniform and log-normal distributions, respectively. In Figure 9 we compare standard and stratified Monte Carlo estimators for all the models above, assuming a computational budget $N=1024$ and training NeurAM using a dataset of size $M=100$. The uniform NeurAM-based stratification is computed using $S=16$ strata, except for $\mathcal{Q}_{1}$ for which we use $S=8$. In the first two test cases, characterized by a relatively low input dimensionality, we also plot the estimator based on a Cartesian stratification on a regular multidimensional grid. Note that a Cartesian grid is not sustainable in high dimensions. In fact, even choosing two strata per dimension, in, e.g., 10 dimensions, this would lead to $10242^{10}=$ 1024 strata in total, saturating the computational budget. We observe that our methodology outperforms standard stratification – assuming this can be computed, as for $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ – and provides effective variance reduction even for high dimensional problems. Regarding the model $\mathcal{Q}_{4}$, we finally remark that the NeurAM surrogate is inaccurate, giving an approximation error of $\sim 23\%$. Nevertheless, the variance reduction is significant, confirming that a highly accurate surrogate model is not essential to improve the estimation even in high dimensions.

### v2

@@ 4.2 Performance of the heuristic algorithm
sMC (u) sMC (h) sMC(o)
Allocation (1) 4.62e-6 (2.29e-02) 3.10e-6 (1.53e-02) 3.11e-6 (1.54e-02)
Allocation (2) 6.26e-6 (3.10e-02) 3.69e-6 (1.83e-02) 3.65e-6 (1.81e-02)
[…]
Figure 5: TOP: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the proposed estimator $\widehat{q}_{\mathrm{sMC}}$ (solid line), for different allocation strategies and stratification approaches: uniform (u), halved (h), optimal (o). The gray dash-dotted vertical line represents the exact value of the quantity of interest. MIDDLE: Mean squared error (MSE) of the estimators to be compared with the value 2.02e-4 for Monte Carlo obtained with $N=1000$. The numbers in parentheses denote the ratio between the MSE of each estimator and the Monte Carlo reference value. BOTTOM: MSE as a function of the computational budget $N$.
We now investigate possible improvements achieved by the heuristic algorithm introduced in Section˜2.3. We fix $M=100$, $K=10^{6}$, $S=10$, assume a computational budget $N=1000$, and test two allocation strategies, as well as two alternatives to iteratively refine the strata. The comparison with standard Monte Carlo and Monte Carlo with Cartesian stratification is shown in Figure˜5. We observe that the heuristic algorithm is always able to outperform uniform stratification, but we do not notice a significant difference between following a bisection approach, i.e., dividing in two equal parts the interval that contributes the most to the final variance, and selecting the optimal new stratum at each step. This behavior can also be seen from the MSE as a function of the computational budget $N=250,500,1000$. Therefore, in order to limit the additional computational overhead, we suggest to use the former approach. We finally remark that, as predicted by theory (see equation (2.20)), the optimal allocation (1) gives a variance which is always smaller than the one produced by a proportional allocation (2).
@@ 4.3 Comparison with active subspaces
[Uncaptioned image] [Uncaptioned image]
[…]
[Uncaptioned image] [Uncaptioned image]
[…]
[Uncaptioned image] [Uncaptioned image]
[…]
$S=2$ $S=8$
sMC (AS) 6.53e-5 (3.31e-01) 2.77e-5 (1.41e-01)
sMC (NeurAM) 6.97e-5 (3.54e-01) 8.42e-6 (4.27e-02)
[…]
Figure 6: TOP: AS and NeurAM-based stratifications of the domain for the model $\mathcal{Q}_{0}$, varying the number of strata $S=2,8$. MIDDLE-TOP: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and stratified Monte Carlo estimator $\widehat{q}_{\mathrm{sMC}}$ (solid line) with both AS and NeurAM-based stratification, varying the number of strata $S=2$ (left) and $S=8$ (right). The gray dash-dotted vertical line represents the exact value of the quantity of interest. MIDDLE-BOTTOM: Mean squared error (MSE) of the estimators to be compared with the value 1.97e-4 for Monte Carlo obtained with $N=1024$. The numbers in parentheses denote the ratio between the MSE of each estimator and the Monte Carlo reference value. BOTTOM: MSE as a function of the computational budget $N$.
NeurAM has the advantages of being nonlinear and entirely data-driven, requiring no knowledge of the model gradient, but it is not the only available option to improve stratified sampling. In this section, we demonstrate that stratified sampling can be combined with other dimensionality reduction techniques. However, we also show that NeurAM outperforms linear state-of-the-art approaches such as the active subspace method [CDW14]. Inspired by [ZGS24b], we build a map from the original domain to the unit interval as follows. Let $\mathcal{G}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a transformation that maps a standard Gaussian $\mathcal{N}(0,I)$ into the input distribution $\mu$, i.e. $\mathcal{G}_{\#}\mathcal{N}(0,I)=\mu$, and define the reparameterized model
@@
where the function $\operatorname{erf}$ is evaluated componentwise, otherwise the map can be computed using, e.g., normalizing flows [KPB21]. Then, let $B$ be the matrix
@@
which is symmetric positive semidefinite and therefore has positive eigenvalues. We denote by $v$ the eigenvector corresponding to the largest eigenvalue, representing the one-dimensional active subspace, and normalize it to have unit norm. We note that the map $\mathcal{G}$ is important because Gaussian distributions are preserved under linear transformations and, in particular, if $z\sim\mathcal{N}(0,I)$ then $v^{\top}z\sim\mathcal{N}(0,1)$. Therefore, a point $x\in\mathbb{D}$ can be mapped into the unit interval $[0,1]$ through the one-dimensional AS, by applying the inverse CDF of the standard Gaussian distribution to $v^{\top}\mathcal{G}^{-1}(x)$. Specifically, using the notation of Section˜2.1, equation (2.12) now reads
@@
In Section˜4.3 we compare the stratification provided by AS with the NeurAM-based stratification for a uniform division of the unit interval into $S=2$ and $S=8$ strata, and a computational budget of $N=1024$. The distribution of the corresponding stratified Monte Carlo estimators with proportional allocation is also shown and compared to that of standard Monte Carlo, which is consistently outperformed. This behavior can also be seen from the MSE as a function of the computational budget $N=64,128,256,512,1024$. While the results are comparable for $S=2$, we observe that the estimator based on nonlinear dimensionality reduction achieves significantly greater variance reduction than linear techniques when the number of strata increases, e.g., $S=8$. This improvement is primarily due to the greater expressiveness of nonlinear methods, which are better able to follow the contour lines of nonlinear models.
@@ 4.4 Stratified multifidelity estimators
sMC sMFMC
Allocation (1) 6.84e-6 (6.77e-02) 3.43e-6 (3.40e-02)
Allocation (2) 9.06e-6 (8.97e-02) 6.32e-6 (6.26e-02)
[…]
Figure 7: TOP: Comparison between standard (multifidelity) Monte Carlo estimators $\widehat{q}_{\mathrm{MC}},\widehat{q}_{\mathrm{MFMC}}$ (dashed line) and the proposed estimators $\widehat{q}_{\mathrm{sMC}},\widehat{q}_{\mathrm{sMFMC}}$ (solid line), for different allocation strategies. The gray dash-dotted vertical line represents the exact value of the quantity of interest. MIDDLE: Mean squared error (MSE) of the estimators to be compared with the value 1.01e-4 for Monte Carlo. The numbers in parentheses denote the ratio between the MSE of each estimator and the Monte Carlo reference value. Standard multifidelity Monte Carlo achieves an MSE of 3.22e-5 (3.19e-01). BOTTOM: MSE as a function of the computational budget $N$.
In this section, we test the performance of estimators combining NeurAM stratification and multifidelity variance reduction, as discussed in Section˜3. Let $\mathcal{Q}^{\mathrm{HF}}=\mathcal{Q}_{0}$, and consider a low-fidelity model of the form
@@
for which we assume a cost ratio $w=0.01$, to reflect cost differences in realistic applications. This poorly correlated low-fidelity model has already been used in [GEG18, ZGS24a, ZGS24b, ZGS25]. Following [ZGS25, Section 3.2], we then train NeurAM for both models using $M=100$, and re-parameterize the low-fidelity model using
@@
resulting in higher correlations with respect to the high-fidelity model. In Figure˜7 we compare (stratified) Monte Carlo estimators with (stratified) multifidelity Monte Carlo estimators for both the allocation strategies, assuming a computational budget $N=2,000$ and using a uniform stratification with $S=5$ strata. First, we observe that stratification is always beneficial, and stratified estimators outperform multifidelity estimators. This behavior can also be seen from the MSE as a function of the computational budget $N=125,250,500,1000,2000$. Then, we notice that leveraging stratification for multifidelity estimators allows us to obtain a significant variance reduction. Finally, we remark that, when using the optimal allocation strategy (1), the improvement with respect to single-fidelity estimators is greater. This is in agreement with Remarks˜3.3 and 3.2, where a weaker sufficient condition is provided for allocation (1) to achieve variance reduction.
4.5 Higher-dimensional models
MC sMC (standard – best) sMC (NeurAM)
$\mathcal{Q}_{1}$ 6.48e-03 (1) 6.31e-03 (9.74e-01) 2.36e-03 (3.65e-01)
$\mathcal{Q}_{2}$ 8.22e-03 (1) 2.91e-03 (3.54e-01) 8.96e-05 (1.09e-02)
$\mathcal{Q}_{3}$ 7.81e-01 (1) 2.87e-01 (3.67e-01) 2.55e-02 (3.26e-02)
$\mathcal{Q}_{4}$ 1.96e-04 (1) 1.12e-04 (5.73e-01) 3.08e-05 (1.57e-01)
[…]
Figure 8: TOP: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the stratified estimators $\widehat{q}_{\mathrm{sMC}}$ (solid line) with both standard – either a regular Cartesian grid ($\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$) or uniform stratification along one dimension ($\mathcal{Q}_{3}$ and $\mathcal{Q}_{4}$) – and NeurAM-based stratification, for the four models in Section˜4.5. MIDDLE: Variance of the estimators obtained with $N=2000$. For the uniform stratification along one dimension of the models $\mathcal{Q}_{3}$ and $\mathcal{Q}_{4}$, the best value is considered. The numbers in parentheses denote the ratio between the variance of each estimator and the Monte Carlo reference value. BOTTOM: Variance as a function of the computational budget $N$.
After investigating the convergence of the proposed approach and its sensitivity to hyperparameters, we now focus on how these properties scale to higher dimensions. Consider the following models:
  • $\mathcal{Q}_{1}\colon\mathbb{R}^{3}\to\mathbb{R}$ proposed in [IsH90] which exhibits strong nonlinearity and nonmonotonicity
@@ 4.5 Higher-dimensional models
  • $\mathcal{Q}_{2}\colon\mathbb{R}^{4}\to\mathbb{R}$ employed in [GCS17] which models the average velocity of a steady, incompressible, and laminar flow of an electrically conducting fluid between two infinite parallel plates in the presence of a uniform magnetic field (see Hartmann problem, e.g. [SPC16])
@@
  • $\mathcal{Q}_{3}\colon\mathbb{R}^{8}\to\mathbb{R}$ proposed in [HaG83] as a model for the flow of water through a borehole
@@
    $\mathcal{Q}_{4}(x)=\prod_{i=1}^{10}\frac{2\left\lvert x_{i}\right\rvert+i}{1+i}.$ (4.12)
@@
[…]
@@
where $\log\mathcal{U}$ and $\log\mathcal{N}$ denote the log-uniform and log-normal distributions, respectively. In Figure˜8 we compare standard and stratified Monte Carlo estimators for all the models above, assuming a computational budget $N=1024$ and training NeurAM using a dataset of size $M=1024$. The uniform NeurAM-based stratification is computed using $S=16$ strata, except for $\mathcal{Q}_{1}$ for which we use $S=8$. In the first two test cases, characterized by a relatively low input dimensionality, we also plot the estimator based on a Cartesian stratification on a regular multidimensional grid. Note that a Cartesian grid is not sustainable in high dimensions. In fact, even choosing two strata per dimension, in, e.g., 10 dimensions, this would lead to $2^{10}=1024$ strata in total, saturating the computational budget. Therefore, for the last two test cases, in Figure 8 we also plot stratified Monte Carlo estimators obtained by partitioning each dimension uniformly, one at a time. We note that the performance in these cases is poor, with results being roughly equivalent across all dimensions. We observe that our methodology outperforms standard stratification and provides effective variance reduction even for high dimensional problems. This behavior can also be seen from the plot of the variance as a function of the computational budget $N=64,128,256,512,1024$. Regarding the model $\mathcal{Q}_{4}$, we finally remark that the NeurAM surrogate is inaccurate, giving an approximation error of $\sim 23\%$. Nevertheless, the variance reduction is significant, confirming that a highly accurate surrogate model is not essential to improve the estimation even in high dimensions.
4.6 Comparison with other variance reduction techniques
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[…]
$d=5$ $d=10$ $d=20$
MC 4.68e-04 (1) 5.05e-04 (1) 5.33e-04 (1)
sMC 1.74e-05 (3.72e-02) 2.88e-05 (5.70e-02) 4.20e-05 (7.89e-02)
LHS 1.43e-04 (3.05e-01) 3.70e-04 (7.32e-01) 4.51e-04 (8.47e-01)
qMC 2.88e-06 (6.15e-03) 1.19e-04 (2.36e-01) 3.77e-04 (7.09e-01)
[…]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[…]
Figure 9: TOP: Comparison of standard Monte Carlo $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the proposed NeurAM-based stratified estimator $\widehat{q}_{\mathrm{sMC}}$ with Latin Hypercube Sampling $\widehat{q}_{\mathrm{LHS}}$ and Quasi-Monte Carlo using Sobol sequences $\widehat{q}_{\mathrm{qMC}}$ (solid line), for different dimensions $d=5,10,20$. The gray dash-dotted vertical line represents the exact value for the expected value of the model’s output. MIDDLE: Mean squared error (MSE) of the estimators obtained with $N=1024$. The numbers in parentheses denote the ratio between the MSE of each estimator and the Monte Carlo reference value. BOTTOM: MSE as a function of the computational budget $N$.
Different sampling-based variance reduction techniques have been proposed in the literature beyond stratified sampling. In this section, we compare our proposed methodology for stratification with two well-established strategies: Latin Hypercube Sampling (LHS) and randomized Quasi-Monte Carlo (qMC). We note that deterministic qMC is not considered here, as our focus is on variance reduction for unbiased estimators and, by construction, deterministic qMC does not possess variance. It is well-known that both LHS and randomized qMC significantly outperform standard MC and achieve higher convergence rates for smooth models in low dimensions. Also, depending on the specific application, LHS and randomized qMC can have different advantages. For instance, with LHS (and stratification) it is possible to generate sets of samples with arbitrary number of points, whereas for qMC (either deterministic or randomized) the convergence properties are guaranteed only if the number of samples corresponds to a power of two, which can be overly restrictive for many applications. Moreover, for both LHS and qMC, the dimensionality of the problem has an effect on their performance. To illustrate the effect of dimensionality, we introduce an analytical model that can be easily scaled in term of its dimension. Consider the model $\mathcal{Q}^{(d)}\colon\mathbb{R}^{d}\to\mathbb{R}$ defined as
$\mathcal{Q}^{(d)}(x)=\sin\left(\sum_{i=1}^{d}x_{i}\right),$ (4.14)
where $d$ denotes the input dimensionality, and assume a uniform input probability distribution $\mu=\mathcal{U}([-1,1]^{d})$, which implies $q_{d}=\operatorname{\mathbb{E}}^{\mu}[\mathcal{Q}_{d}(X)]=0$. We then set $M=1000$ and $K=10^{6}$ to learn the NeurAM and the CDF, respectively. In Section˜4.6, for different values of the dimension $d=5,10,20$, we compare independent repetitions of our stratified estimator with $S=16$ strata to the same number of independent repetitions for standard Monte Carlo, LHS, and randomized qMC, assuming a computational budget $N=1024$. Both LHS and qMC are implemented using the scipy.stats.qmc Python package. We also note that the randomized version of qMC Sobol’ sequences in scipy is based on digital shift and includes linear matrix scrambling. Similar results can be obtained using randomized digital nets from the QMCPy Python package, whereas randomized Halton sequences produce results with higher variance. For all estimators, in Section˜4.6 we also report the mean squared error (MSE) for $N=1024$ and plot it as a function of an increasing computational budget $N=64,128,256,512,1024$ (we limit the computational budget to powers of two to avoid penalizing qMC). While LHS performs overall only slightly better than standard Monte Carlo, randomized qMC provides a substantial improvement over both Monte Carlo and our stratified estimator in terms of MSE and convergence rate for $d=5$. Similar results can also be observed for the functions used in the previous sections due to their dimensionality. However, qMC variance reduction compared to MC degrades as the dimension increases to $d=10$ and offers no improvement for $d=20$. In contrast, the NeurAM-based stratified Monte Carlo estimator is not significantly affected by increasing dimensionality, resulting in a more robust estimator that consistently reduces the variance of standard Monte Carlo across all three considered dimensions. Finally, we also remark that the evaluation of the estimator variance, which is needed for the construction of a confidence interval, is straightforward for Monte Carlo and stratified estimators (see, e.g., equation Eq.˜2.15 for our stratified estimator), but it requires repetitions for qMC, which reduces the effective computational budget available for estimating the statistics of interest.
4.7 Darcy flow
[Uncaptioned image] [Uncaptioned image]
[…]
Figure 10: TOP: A single realization of the permeability $\mathcal{K}$, $p$, $v_{x}$, $v_{y}$ for the Darcy flow problem. BOTTOM-LEFT: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and NeurAM-based stratified estimator $\widehat{q}_{\mathrm{sMC}}$. BOTTOM-RIGHT: Variance as a function of the computational budget $N$.
In the last numerical experiment, we consider the two-dimensional Darcy flow problem, which models a single phase, steady-state flow through a random permeability field. Let $\Omega=[0,1]^{2}$ be the spatial domain representing an idealized oil reservoir, and let $p\colon\Omega\to\mathbb{R}$ and $v\colon\Omega\to\mathbb{R}^{2}$ be the pressure and velocity field of the fluid, respectively, which satisfy the equations
$-\nabla\cdot(\mathcal{K}\nabla p)$ $=f,$ $\text{in }\Omega,$ (4.15)
$v$ $=-\mathcal{K}\nabla p,$ $\text{in }\Omega,$
where $\mathcal{K}\colon\Omega\to\mathbb{R}_{+}$ is the permeability field and $f\colon\Omega\to\mathbb{R}$ is a source term. Moreover, let $\Gamma_{D}^{l}=\{0\}\times[0,1]$, $\Gamma_{D}^{r}=\{1\}\times[0,1]$, $\Gamma_{N}=\partial\Omega\setminus(\Gamma_{D}^{l}\cup\Gamma_{D}^{r})$, and impose the following Dirichlet and Neumann boundary conditions
$p$ $=1,$ $\text{on }\Gamma_{D}^{l},$ (4.16)
$p$ $=0,$ $\text{on }\Gamma_{D}^{r},$
$\nabla p\cdot\mathrm{n}$ $=0,$ $\text{on }\Gamma_{N}.$
The random permeability $\mathcal{K}$ is modeled as $\mathcal{K}=e^{G}$, where $G$ is a Gaussian random field $G\sim\mathcal{N}(0,\Xi)$ with Gaussian covariance function
$\Xi((x_{1},y_{1}),(x_{2},y_{2}))=\tau^{2}\exp\left[-\frac{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}{\ell^{2}}\right].$ (4.17)
In order to obtain finite-dimensional samples of $G$, we employ the truncated Karhunen–-Loève expansion with $\kappa$ terms
$G(x,y)=\sum_{j=1}^{\kappa}\xi_{j}\sqrt{\theta_{j}}\psi_{j}(x,y),$ (4.18)
where $(\psi_{j})_{j=1}^{\infty}$ are eigenfunctions of the covariance operator $\Xi$ and form an orthonormal basis of $L^{2}(\Omega)$, $(\theta_{j})_{j=1}^{\infty}$ are the corresponding eigenvalues, and $(\xi_{j})_{j=1}^{\kappa}$ are independent and identically distributed standard Gaussian random variables, i.e., $\xi_{j}\sim\mathcal{N}(0,1)$. Our goal is then estimating the expectation of the quantity of interest $\mathcal{V}$ given by
$\mathcal{V}=\int_{0}^{1}\int_{0}^{1}(v_{x}(x,y)^{2}+v_{y}(x,y)^{2})\,\mathrm{d}x\,\mathrm{d}y,$ (4.19)
where $v_{x}$ and $v_{y}$ denote the two components of the velocity field $v$.
The Darcy flow simulations are performed using the FEniCS computing platform [LMW12] with a spatial discretization of $64$ elements per dimension. A representative realization of the permeability field $\mathcal{K}$, the pressure $p$, and the velocity components $v_{x}$ and $v_{z}$ is shown in Section˜4.7. In this experiment, we set the source term to zero, $f=0$, choose a truncation level $\kappa=100$, and fix the standard deviation and correlation length of the covariance function to $\tau=0.2$ and $\ell=0.5$, respectively. The lower part of the same figure compares the performance of the NeurAM-based stratified estimator with that of a standard Monte Carlo estimator. We employ $M=2000$ samples to train the NeurAM, $K=10^{6}$ for the CDF, and assume a total computational budget of $N=512$. Unlike the previous test cases, due to the increased cost of each simulation, we employ 20 independent evaluations of each estimator for variance estimation. The resulting variances are $\operatorname{Var}[\widehat{q}_{\mathrm{MC}}]=\text{1.40e-4}$ and $\operatorname{Var}[\widehat{q}_{\mathrm{sMC}}]=\text{4.20e-6}$, yielding a variance ratio of 3.00e-2. In addition, we report the estimator variance as a function of the computational budget $N=32,64,128,256,512$. These results, obtained for random inputs in $d=\kappa=100$ dimensions, show that the proposed stratification approach significantly reduces Monte Carlo variance even in high-dimensional settings.